A K-56 atom (potassium-56) has a nucleus with 19 protons and an atomic mass of 56.0085186 u. What is the nuclear binding energy per nucleon in this atom?

Group of answer choices

429 MeV

22.6 MeV

0.461 MeV

7.66 MeV

11.6 M
Answer #1

429.4644 Mev so!- K-5672=18 N = 56-19 - 37 The Binding energy is EB = (zmu tumn-mk-sq) c2 >> Ep = (1931.007840 + 37x1.008665